### v1

@@ Abstract
To determine the causal effect of a treatment using observational data, it is important to balance the covariate distributions between treated and control groups. However, achieving balance can be difficult when treated and control groups lack overlap. In the presence of limited overlap, researchers typically choose between two types of methods: 1) methods (e.g., inverse propensity score weighting) that imply traditional estimands (e.g., ATE) but whose estimators are at risk of variance inflation and considerable statistical bias; and 2) methods (e.g., overlap weighting) which imply a different estimand, thereby changing the target population to reduce variance. In this work, we introduce a framework for characterizing estimands by their target populations and the statistical performance of their estimators. We introduce a bias decomposition that encapsulates bias due to 1) the statistical bias of the estimator; and 2) estimand mismatch, i.e., deviation from the population of interest. We propose a design-based estimand selection procedure that helps navigate the tradeoff between these two sources of bias and variance of the resulting estimators. Our procedure allows the analyst to incorporate their domain-specific preference for preservation of the original population versus reduction of statistical bias. We demonstrate how to select an estimand based on these preferences by applying our framework to a right heart catheterization study.
@@ Introduction
A primary goal of healthcare research is to determine the causal effect of a policy or treatment on health outcomes. Assessing causal effects requires an unconfounded comparison of treated and control groups such that the two groups have similar, or balanced, characteristics. In cases where it is not possible or ethical to balance groups through randomization, observational data must be used to assess the effectiveness of treatment. In observational data, an individuals’ covariates often influence the treatment decision which results in imbalance of the treated and control covariate distributions and, without adjustment, treatment effect estimates with the potential for considerable bias. Recent research has produced a wide variety of methods to adjust for such an imbalance. However, there can be a lack of overlap in some parts of the distribution of covariates, making it impossible to fully mitigate imbalance, resulting in large estimator bias and variance. Alternate methods aim to reduce estimator variance, through either achieving only approximate balance or by changing the population of interest such that the new population has sufficient overlap to balance the treated and control groups. Yet, these approaches, respectively, increase estimator bias or shift the target estimand away from the originally-targeted estimand. In this paper, we develop a framework for navigating the tradeoffs between preserving the original estimand while minimizing the bias and variance that can result from imbalance.
@@
There are many methods for balancing covariates between groups (e.g., regression, matching, weighting), however, we will focus on weighting methods. Weighting methods reweight the data in order to achieve covariate balance between the treated and control groups. One set of such methods seek to directly balance the covariates by deriving weights subject to covariate constraints using optimization procedures (Hainmueller, 2012; Zubizarreta, 2015; Wang and Zubizarreta, 2020; Huling and Mak, 2024). Another set of methods rely on inverse-probability weighting (IPW), where each individual in the data is inverse weighted by the sample probability of receiving their assigned treatment (Rosenbaum and Rubin, 1983; Hahn, 1998; Robins, Hernán, and Brumback, 2000; Hirano and Imbens, 2001; Hirano, Imbens, and Ridder, 2003; Imbens, 2004). These methods are model-based as the propensity score, the probability of being assigned treatment, must be estimated from observed covariates to derive the weights. Both types of weights can be utilized to target different estimands defined over a specific population; for example, standard IPW weights correspond to the average treatment effect (ATE) which is defined over the total sample population. Other estimands that are commonly targeted are the average treatment effect on the treated (ATT) and on the control (ATC).
Weighting methods are popular and imply estimands with meaningful populations. Yet, when the treated and control covariate distributions lack overlap, achieving balance through these methods can be unattainable, resulting in estimator bias. When there is a lack of overlap, the use of either type of weighting methods can result in very large weights for some observations; large weights yield estimators with inflated variance and the potential for considerable bias (Busso, DiNardo, and McCrary, 2014; Hong, Leung, and Li, 2020). Direct balancing methods often provide more stable and smaller weights than propensity score methods; however, using these methods when there is a lack of overlap can still result in estimators with large variance, and in some cases, no solutions to the optimization problem (Hainmueller, 2012; Chattopadhyay, Hase, and Zubizarreta, 2020). Similar issues arise with matching methods, where large differences between the treated and control groups can result in potential bias in the treatment effect estimator due to imbalance. Thus, employing these methods for covariate balancing when there is lack of overlap produces an estimand with an interpretable and/or meaningful population of interest but whose estimator may have substantial bias and variance.
@@
A variety of new weighting methods have been created to mitigate the impacts of lack of overlap on causal estimators. For the majority of these methods, reducing the variance of the estimator is the primary goal. Within this set of methods there are two types: 1) estimator variance is reduced through increasing estimator bias; and 2) estimator variance is reduced through changing the estimand of interest such that the corresponding population has better overlap. Within this first type of method, Wang and Zubizarreta (2020) explore a class of direct balancing weights called minimal weights that aim to minimize variance when there is a lack of overlap. Minimal weights provide approximate balance, rather than exact balance of the covariate distributions, which tends to increases the bias of the estimator while reducing the variance. While estimator bias may be increased with this method, the estimand and therefore the target population remains unchanged.
@@
Most new methods fall into the second category of methods that modify the estimand, and consequently the target population, either explicitly or implicitly. Crump et al. (2009) formalize modifying the estimand through propensity score trimming while Li and Greene (2013) and Yang and Ding (2018) propose continuous weights that approximate pair matching and propensity score trimming, respectively One of the more popular methods is overlap weights (Li, Morgan, and Zaslavsky, 2018; Li, Thomas, and Li, 2019) (with a corresponding estimand of the ATO), which weight each individual by the probability of being assigned to the opposite group. Overlap weights emphasize individuals with propensity scores around $0.5$, i.e. those with so-called clinical equipoise. When the variance of outcomes is constant, overlap weights minimize the asymptotic variance of the weighted average treatment effect estimator across all possible estimands characterized by weighted populations. Simulation studies have shown that overlap weights perform well under model misspecification and can improve statistical power compared to IPW weights (Zhou, Matsouaka, and Thomas, 2020; Mao, Li, and Greene, 2019). Of note, matching methods that address lack of overlap also change the population of interest by identifying an optimal matched population; however, the primary goal of these methods is to minimize estimator bias, rather than variance, through achieving balance (Rosenbaum, 2012; de los Angeles Resa and Zubizarreta, 2016; Visconti and Zubizarreta, 2018). For all these methods, the estimand of interest, and therefore the population, is changed in order to achieve better statistical performance of the estimator.
@@
There are a variety of scenarios in which it is reasonable to deviate from estimands that aggregate over the entire sample population; for example, some sample populations are a convenience sample rather than a meaningful population (Rosenbaum, 2012) and as such, the ATE over such a population may not have clinical relevance. However, in many other scenarios the sample population or treated population is meaningful, making it still desirable to estimate the ATE or ATT even in the presence of low overlap. Li, Morgan, and Zaslavsky (2018) and Mao, Li, and Greene (2019) both argue that overlap weights target a population of interest; people who have propensity scores around $0.5$ may be the individuals for whom clinical treatment guidelines are unclear. While this is often true, in practice treatments may not be assigned equitably along demographic factors (Faigle et al. (2017); Raeisi-Giglou et al. (2022); Tchikrizov et al. (2023)); in these cases, it is unclear whether individuals at equipoise in the data truly reflect a group of people for whom clinical guidelines are unknown. Regardless of whether clinical equipoise is of interest, it is also often hard to describe the exact population that is induced by overlap weights. Within methods that address lack of overlap, there has been limited development on constructing an estimand that also has an interpretable population. Cardinality matching provides an option for constraining the matched set to be similar to some population of interest (Visconti and Zubizarreta, 2018). Traskin and Small (2011) and Fogarty et al. (2016) both propose methods for selecting an interpretable population with better overlap, through tree and discrete optimization techniques respectively, though both methods can still produce estimators with high variance.
@@
There are clear tradeoffs between retaining an estimand with a meaningful target population, estimator bias, and variance among current methods for treatment effect estimation when there is a lack of overlap. Standard weighting methods target estimands with meaningful populations, however there is risk of estimator bias and inflated variance. Methods such as minimal weights and overlap weights aim to either derive a set of weights or identify an estimand that minimizes estimator variance, but in doing so they sacrifice estimator bias or eschew the original target population. Of note, most methods described above aim to improve or maintain only one of estimator bias, variance, or the target population; there are few, if any, tools for the quantifying the extent with which the other two characteristics are influenced. Different applications may prioritize these three characteristics differently; in some cases statistical performance may be the primary priority, while in others it may be important to maintain the population of interest. However, as far as we know, there are few or no methods that provide tools to flexibly explore and quantify the tradeoffs between all three characteristics.
@@
In this work, we present a conceptual framework and design-based procedure for characterizing and selecting an estimand according to its target population, estimator bias, and variance, allowing the analyst to explicitly prioritize any of these properties. As a part of the conceptual framework, we introduce a bias decomposition that encapsulates bias due to 1) the statistical bias of the estimator; and 2) estimand mismatch i.e., deviation from the population of interest. We present a set of estimands from which to select an optimal estimand, where the ATE, ATT, ATC, and ATO are special cases of the set. To characterize this set of estimands, we propose two design-based metrics that target each part of the bias decomposition. Furthermore, we propose an estimand selection procedure that clearly illustrates the bias-variance tradeoffs of potential estimands. We analyze synthetic data from a variety of data generating scenarios and data from a right heart catheterization study to demonstrate the effectiveness of our method in navigating these tradeoffs. We present all methods and results in reference to the ATE population as the population of interest, however, all methods can apply to other estimands/populations such as the ATT or ATC.
@@
The remainder of the paper is outlined as follows. In Section 2 we introduce notation and our conceptual framework for characterizing and selecting estimands. Section 3 describes the design-based metrics for characterizing both types bias and the proposed estimand selection procedure. Section 4 validates our metrics and evaluates the performance of our estimation selection procedure in comparison to the ATE and ATO through simulation studies. In Section 5 we apply our entire procedure to reanalyze a study on right heart catheterization. Finally, Section 6 discusses the results and future work.
@@
Consider a sample $\{(Z_{i},Y_{i},\bm{X}_{i})\}_{i=1}^{n}$ of size $n$ from a population. For individual or unit $i$, let $Z_{i}=z$, indicate belonging to the treatment ($z=1$) or control group ($z=0$), $Y_{i}$ be the outcome, and $\bm{X}_{i}=(X_{i1},\ldots X_{ip})$ be a $p$ length vector of covariates. Let $e(\bm{x})=\Pr(Z_{i}=1|\bm{X}_{i}=\bm{x})$ be the propensity score. We use the potential outcomes framework (Neyman (1990); Rubin (1974, 1978); Hernan and Robins (2024)), where $Y(z)$ is the outcome that would be observed if assigned to treatment group $z$. In practice, only one potential outcome is observed per individual. We assume the standard stable unit treatment value assumption (SUTVA) that states the potential outcomes for each individual are unaffected by the treatment assignments of other individuals and that there is only one version of treatment, which implies that $Y_{i}=Y_{i}(Z_{i})$. We also assume that the treatment effect is strongly unconfounded such that $\{Y(0),Y(1)\}{\,\perp\!\!\!\perp\,}Z|\bm{X}$, which requires that there are no unmeasured confounders. Finally, we assume positivity (Rosenbaum and Rubin (1983)), which requres that $10<e(\bm{x})<$ 1 such that all individuals have a non-zero probability of being assigned to either the treatment or control group. Given these assumptions and using potential outcomes notation, the conditional average treatment effect (CATE) is $\tau(\bm{x})=\mu_{1}(\bm{x})-\mu_{0}(\bm{x})=E\{Y(1)-Y(0)|\bm{X}=\bm{x}\}$. Often, we desire to compute the ATE not with respect to a single $\bm{x}$ but over a specific target population or covariate distribution. Let $f(\bm{x})$ be the marginal distribution of the covariates and $h(\bm{x})$ some function such that $f(\bm{x})h(\bm{x})$ is the density of a target population. Then the estimand, $\tau_{h}$, averages the CATE over the target population,
@@
Note that each unique $h(\bm{x})$ defines a different target population and thus a different causal estimand, $\tau_{h}$. Throughout, we assume that $\tau(\bm{x})$ can vary with $\bm{x}$, such that averaging over different populations (i.e., different causal estimands) may yield different results. For a given weight function $h(\bm{x})$, Li, Morgan, and Zaslavsky (2018) propose a class of propensity score balancing weights such that weighted covariate distributions are balanced between the two treatment groups,
$w_{1}(\bm{x})$ $=\frac{h(\bm{x})}{e(\bm{x})}\>\>\>\text{for}\>\>\>Z=1,\text{ and}$ (1)
$w_{0}(\bm{x})$ $=\frac{h(\bm{x})}{1-e(\bm{x})}\>\>\>\text{for}\>\>\>Z=0.$ (2)

### v2

@@ Abstract
Estimating the causal effect of a treatment or health policy with observational data can be challenging due to an imbalance of and a lack of overlap between treated and control covariate distributions. In the presence of limited overlap, researchers choose between 1) methods (e.g., inverse probability weighting) that imply traditional estimands but whose estimators are at risk of considerable bias and variance; and 2) methods (e.g., overlap weighting) which imply a different estimand, thereby modifying the target population to reduce variance. We propose a framework for navigating the tradeoffs between variance and bias due to imbalance and lack of overlap and the targeting of the estimand of scientific interest. We introduce a bias decomposition that encapsulates bias due to 1) the statistical bias of the estimator; and 2) estimand mismatch, i.e., deviation from the population of interest. We propose two design-based metrics and an estimand selection procedure that help illustrate the tradeoffs between these sources of bias and variance of the resulting estimators. Our procedure allows analysts to incorporate their domain-specific preference for preservation of the original research population versus reduction of statistical bias. We demonstrate how to select an estimand based on these preferences with an application to right heart catheterization data.
@@ Introduction
A primary goal of healthcare research is to determine the causal effect of a treatment or policy on health outcomes. Estimating causal effects using observational data requires 1) an unconfounded comparison of treated and control groups such that the two groups have similar, or balanced, characteristics; and 2) overlap between the treated and control groups such that all observations have a non-zero probability of being assigned either treatment or control. While there are a wide variety of methods to adjust for covariate imbalance, overlap is less well-understood and there are relatively fewer methods that address its impacts on the estimation of causal effects. Some methods aim to reduce variance inflation caused by limited overlap through either achieving only approximate balance or by modifying the population of interest to a new population has sufficient overlap. Yet, these approaches, respectively, increase estimator bias or shift the target estimand away from the original estimand of scientific interest. It is also often unclear which method, or which estimand, is best suited for a given scientific question and degree of overlap. In this paper, we develop a framework for simultaneously navigating the tradeoffs between preserving the original target population and minimizing the bias and variance that can result from imbalance and lack of overlap.
@@
There are many methods that balance the covariates between groups (e.g., regression, matching, weighting). We focus on weighting methods which reweight the data in order to achieve covariate balance. One set of such methods seek to directly balance the covariates by deriving weights subject to covariate constraints using optimization procedures (Hainmueller, 2012; Zubizarreta, 2015; Wang and Zubizarreta, 2020). Another set of methods rely on inverse probability weighting (IPW), where observations are inverse weighted by the sample probability of receiving their assigned treatment (Rosenbaum and Rubin, 1983a; Hahn, 1998; Robins et al., 2000; Hirano and Imbens, 2001; Hirano et al., 2003; Imbens, 2004). Both types of methods can target estimands corresponding to a specific population, such as the average treatment effect defined over the entire (ATE), treated (ATT), or control (ATC) sample population. While these methods target estimands with a meaningful population, a lack of overlap between the treated and control covariate distributions can create very large weights that produce estimators with inflated variance and the potential for considerable bias (Busso et al., 2014; Hong et al., 2020). While direct balancing methods generally provide more stable and smaller weights, a lack of overlap can yield no solutions to the optimization problem (Hainmueller, 2012; Chattopadhyay et al., 2020). Similar issues arise with matching methods, where large differences between the treated and control groups can result in bias in the treatment effect estimator due to imbalance.
@@
A variety of new weighting methods have been created that mitigate the impacts of lack of overlap on estimation. For the majority of these methods, reducing variance of the estimator is the primary goal. Wang and Zubizarreta (2020) explore a class of direct balancing weights called minimal weights which provide approximate, rather than exact balance of the covariate distributions, resulting in estimators with reduced variance but potentially increased bias. Other approaches work by modifying the estimand, and consequently the target population, either explicitly or implicitly to achieve better overlap. Crump et al. (2009) formalize modifying the estimand through propensity score trimming while Li and Greene (2013) and Yang and Ding (2018) propose continuous weights that approximate pair matching and propensity score trimming, respectively. One of the more widely-used such methods is overlap weights (Li et al., 2018, 2019), which minimize the asymptotic variance of the weighted average treatment effect estimator across all estimands characterized by weighted populations under the assumption of homoscedasticity. Overlap weights emphasize individuals with propensity scores around $0.5$, i.e. those with so-called clinical equipoise, targeting the “overlap” (ATO) population. Matching methods also often address lack of overlap by modifying the matched population and frequently operate by discarding hard-to-match samples, improving balance by modifying the population (Rosenbaum, 2012; de los Angeles Resa and Zubizarreta, 2016; Visconti and Zubizarreta, 2018).
@@
There are a variety of scenarios in which it is reasonable to deviate from estimands that aggregate over the entire sample population; for example, some sample populations are a convenience sample (Rosenbaum, 2012) and as such, the ATE may not have clinical relevance. Li et al. (2018) and Mao et al. (2019) both argue that overlap weights target a population of interest; people with propensity scores around $0.5$ may be the individuals for whom clinical treatment guidelines are unclear. While this is often true, in practice treatments and policies may not be assigned equitably along demographic factors (Faigle et al. (2017); Raeisi-Giglou et al. (2022); Tchikrizov et al. (2023)); in these cases, the interpretation of individuals at equipoise is unclear. Regardless of whether clinical equipoise is of interest, it can be hard to describe the exact population that is induced by overlap weights. Therefore, in many scenarios it is still desirable to estimate the ATE or ATT even in the presence of low overlap to preserve a meaningful and/or interpretable population. Among methods that address lack of overlap, there has been limited development on constructing an estimand that also has an interpretable population. Traskin and Small (2011) and Fogarty et al. (2016) propose methods for selecting an interpretable population with better overlap, though both methods can still produce estimators with considerable bias and/or variance.
@@
There are clear tradeoffs between retaining an estimand with a meaningful target population, estimator bias, and variance among current methods for treatment effect estimation when there is a lack of overlap. Standard weighting methods target estimands with meaningful populations, however there is risk of estimator bias and inflated variance. Methods such as minimal weights and overlap weights aim to minimizes estimator variance, but in doing so they either sacrifice estimator bias or eschew the original target population. Most methods described above aim to improve or maintain only one of estimator bias, variance, or the target population, but the extent to which the other two characteristics are influenced for a given analysis is often unclear. Different biomedical or healthcare application may seek to prioritize these three characteristics differently; in some cases statistical performance may be the primary priority, while in others it may be important to maintain the population of interest. However, as far as we know, there are few or no methods that provide tools to flexibly explore and quantify the tradeoffs between all three characteristics.
@@
In this work, we present a conceptual framework and corresponding procedure for characterizing and selecting an estimand according to its target population, estimator bias, and variance, allowing the analyst to explicitly prioritize any of these properties. As part of the conceptual framework, we introduce a bias decomposition that encapsulates bias due to 1) the statistical bias of the estimator; and 2) estimand mismatch i.e., deviation from the population of interest. To leverage this decomposition for estimand selection, we propose two design-based metrics that target each part of the bias decomposition. Furthermore, we propose an estimand selection procedure that enumerates a sequence of estimands corresponding to different tradeoffs between preservation of the target estimand and statistical performance in terms of bias and variance, where the ATE, ATT, ATC, and ATO are special cases of the set of possible estimands under consideration. We analyze synthetic data from a variety of data generating scenarios and data from a right heart catheterization study to demonstrate the effectiveness of our method in navigating these tradeoffs. We present all methods and results in reference to the ATE as the estimand of scientific interest, however, all methods can apply to other estimands/populations such as the ATT.
@@
The remainder of the paper is outlined as follows. In Section 2, we introduce notation and our conceptual framework for characterizing and selecting estimands. Section 3 describes the design-based metrics for characterizing both sources bias and the proposed estimand selection procedure. Section 4 validates our metrics and evaluates the performance of our estimation selection procedure in comparison to the ATE and ATO through simulation studies. In Section 5 we apply our entire procedure to reanalyze a study on right heart catheterization. Finally, Section 6 discusses the results and future work.
@@
$w_{1}(\bm{x})$ $=\frac{h(\bm{x})}{e(\bm{x})}\>\>\>\text{for}\>\>\>Z=1,\text{ and }\>\>\>w_{0}(\bm{x})=\frac{h(\bm{x})}{1-e(\bm{x})}\>\>\>\text{for}\>\>\>Z=0.$ (1)